# Faithful global convergence for the rescaled Consensus–Based Optimization

### v1

@@ Abstract.
We analyze the Consensus-Based Optimization (CBO) algorithm with a consensus point rescaled by a small fixed parameter $\kappa\in(0,1)$. Under minimal assumptions on the objective function and the initial data, we establish its unconditional convergence to the global minimizer. Our results hold in the asymptotic regime where both the time–horizon $t\to\infty$ and the inverse–temperature $\alpha\to\infty$, providing a rigorous theoretical foundation for the algorithm’s global convergence. Furthermore, our findings extend to the case of multiple and non–discrete set of minimizers.
@@ 1. Introduction
Consensus-Based Optimization (CBO) is a derivative–free algorithm designed to tackle large–scale, highly non–convex optimization problems. This algorithm and its analytical foundations were established in [pinnau2017consensus, carrillo2018analytical]. It has since then been extensively studied, both theoretically and practically, demonstrating its effectiveness in solving complex optimization challenges that arise in real–world applications.
Inspired by the collective behavior observed in social and biological systems, CBO exploits the interactions of multiple agents, or particles, to explore the search space in a coordinated and efficient manner. The method uses a stochastic consensus mechanism, where particles are dynamically guided toward regions of lower objective values, ultimately converging to an optimal solution. This decentralized and adaptive search strategy makes CBO particularly well-suited for problems where gradient information is unavailable or unreliable. Motivated by its broad applicability, researchers have extended and refined the original CBO framework to accommodate various settings and demonstrated strong empirical success in solving wide range of global optimization problems [fornasier2021JMLR, ha2022stochastic, borghi2022consensus, herty2024multiscale, carrillo2022consensus, carrillo2021consensus, fornasier2022anisotropic, fornasier2020consensus, fornasier2024truncated, byeon2025consensus, chen2022consensus, wei2025consensus, bonandin2024consensus, cipriani2022zero]. These advancements have further enhanced the algorithm’s robustness and efficiency, making it a versatile tool across a wide range of scientific and engineering disciplines.
@@
Find $\text{Find }\;x_{*}\in\operatorname*{argmin}_{x\in{\mathbb{R}}^{d}}f(x),\,$
@@
where $\sigma>0$ is a real constant, $\rho_{t}^{N}:=\frac{1}{N}\sum_{i=1}^{N}\delta_{X_{t}^{i}}$ is the empirical measure associated to the particle system, and $\{B_{.}^{i}\}_{i=1}^{N}$ are $N$ independent $d$-dimensional Brownian motions. Here, we employ the anisotropic diffusion in the sense that $D(X):=\mbox{diag}(|X_{1}|,\dots,|X_{d}|)$ for any $X\in{\mathbb{R}}^{d}$, which has been proven to handle high-dimensional problems more effectively [carrillo2021consensus, fornasier2022anisotropic]. The current global consensus point $\mathfrak{m}_{\alpha}(\rho_{t}^{N})$ is meant to approximate the global minimum, and is defined with a weighted average of the particles as
@@
This choice of weight function $\omega_{\alpha}^{f}(\cdot)$ is motivated by the well-known Laplace’s principle [miller2006applied, MR2571413], see also [huang2024consensus, Appendix A.2]. Moreover, the particle system is initialized with independent and identically distributed (i.i.d.) data $\{X_{0}^{i}\}_{i=1}^{N}$, where each $X_{0}^{i}$ is distributed according to a given measure $\rho_{0}\in\mathscr{P}({\mathbb{R}}^{d})$ for all $i\in[N]$.
The convergence proof of the CBO method is typically conducted in the context of large-particle limit [gerber2023mean, huang2022mean]. Specifically, rather than analyzing the $N$–particle system (1.1) directly, one considers the limit as $N$ approaches infinity and examines the corresponding McKean–Vlasov process ${\overline{X}}_{\cdot}$, which is governed by the following equation:
@@
$\text{with }\;\rho_{t}^{\alpha}\,:=\,\text{Law}({\overline{X}}_{t}).$
@@ 1.1. The convergence problem in the literature
There are numerous results in the literature on the convergence analysis of CBO methods, which can be broadly categorized into two different approaches. For instance, in [fornasier2024consensus], the authors employ $\mathbb{E}[|\overline{X}_{t}-x_{*}|^{2}]$ as a Lyapunov functional for (1.3). Their convergence result states that for any given accuracy $\varepsilon>0$, there exists a time $T_{*}>0$ depending on $\varepsilon$, such that for $\alpha>\alpha_{0}$ (also depending on $\varepsilon$), the following holds:
@@
This approach has been further extended to establish the convergence of various CBO variants, including those incorporating general constraints [borghi2023constrained, beddrich2024constrained], multi-player games [chenchene2023consensus], min-max problems [borghi2024particle], multiple-minimizer problems [fornasier2025pde], clustered federated learning [carrillo2023fedcbo], and bi-level optimization [trillos2024cb].
However, a limitation of this method is that it does not allow for taking the limits $t\to\infty$ or $\alpha\to\infty$, meaning that convergence is guaranteed only within a neighborhood of $x_{*}$ with accuracy $\varepsilon$.
One possible approach to achieving convergence as $t\to\infty$ is through a variance-based method described in [carrillo2018analytical]. This approach involves defining the variance function
@@
The proof is then divided into two steps. First, it is shown that under well-prepared initial data and appropriately chosen parameters, for any fixed $\alpha$, the function $V(t)$ converges to zero exponentially fast as $t\to\infty$. This implies that $\overline{X}_{t}$ converges to some consensus point $x^{\alpha}$ as $t\to\infty$. Second, it is verified that the consensus point $x^{\alpha}$ can be made arbitrarily close (up to any given accuracy $\varepsilon$) to the global minimizer $x_{*}$ by choosing some fixed $\alpha$ that is sufficiently large. This argument has been further extended to establish the convergence of such as particle swarm optimization [huang2023global], CBO constrained on the sphere [fornasier2021JMLR], CBO for saddle point problems [huang2024consensus], and CBO with mirror descent [bungert2025mirrorcbo]. A similar idea is also employed in the convergence proof at the particle level [ko2022convergence].
However, similar to the first method, this approach does not allow taking the limit $\alpha\to\infty$ and it only guarantees convergence within an $\varepsilon$–neighborhood of the global minimizer.
@@
1.2. A key observation and the rescaled CBO
As it has been observed in [huang2024self], any Dirac measure is invariant for the dynamics (1.3) which makes it difficult to study its asymptotic properties. This non–uniqueness of an invariant probability measure is mainly due to two reasons: in the drift, the term $-X$ is not strong enough with respect to the mean–field term $\mathfrak{m}_{\alpha}(\cdot)$, and the diffusion degenerates. This is the main reason behind our modification of (1.3) whose aim is to address exactly these two mentioned issues, and which has already been proposed in [huang2024uniform, huang2024self, herty2024multiscale]. Hence, we consider a rescaled CBO given by
@@ 1.2. A key observation and the rescaled CBO
$\text{with }\;\rho^{\alpha}_{t}\,:=\,\text{Law}({\overline{X}}_{t}).$
complemented with an initial condition ${\overline{X}}_{0}\sim\rho_{0}\in\mathscr{P}_{4}({\mathbb{R}}^{d})$, where $\mathds{I}_{d}$ is the $d$-dimensional identity matrix, $10<\kappa<$ 1 is a small positive constant, and $\delta>0$ is any positive constant. The name is justified by the fact that the contribution of the mean–field term $\mathfrak{m}_{\alpha}(\cdot)$ is re-scaled with a parameter $\kappa$ when compared to the contribution of $-{\overline{X}}$.
Introducing the parameter $\kappa$ makes all Dirac measures not invariant anymore, and intuitively, creates a disconnection between the limits in the time variable $t\to\infty$ and the inverse–temperature $\alpha\to\infty$, hence allowing to easily handle the convergence of the dynamics towards the (to–be) global minimum.
On the other hand, the positive parameter $\delta\,$ guarantees ellipticity (non–degeneracy) of the PDE satisfied by the invariant probability measure of the dynamics, and grants it nice properties. More precisely, we need $\delta>0$ in order to guarantee existence of an invariant probability measure as in [huang2024self, Proposition 3.4]. Uniqueness then follows using the empirical approximation in [huang2024self, Corollary 3.8] which holds with no positivity requirement on $\delta$. Its value however does not influence the asymptotic analysis that we shall perform in the sequel, or the results that we aim for.
Clearly, the underlying system of $N$ interacting particles is of the form
@@
and $\rho_{t}^{N}:=\frac{1}{N}\sum_{i=1}^{N}\delta_{X_{t}^{i}}$ is the empirical measure associated to the particle system. In the following Figure 1, using standard Euler–Maruyama scheme, we apply our rescaled CBO particle system (1.5) as described in [huang2024uniform, Algorithm 1] to the Rastrigin function $R(x):=10+(x-1)^{2}-10\cos(2\pi(x-1))\,,$ which is a popular benchmark nonconvex function in optimization with multiple local minimizers and a unique global minimizer.
Figure 1. We apply the rescaled CBO particle system (1.5) to the Rastrigin function $R(x)$, which has a unique global minimizer $x_{*}=1$ (the red star). The initial particles (the blue dots) are sampled uniformly in $[2,5]$. The simulation parameters are $N=2*10^{6},\lambda=1,\sigma=0.2,\alpha=10^{15},\delta=1,{\rm d}t=0.01,T=100$ and $\kappa=0.01$. The final output is $\frac{1}{\kappa}\frac{1}{N}\sum_{i\in[N]}X_{T}^{i}\approx\frac{1}{\kappa}{\mathbb{E}}[{\overline{X}}_{T}]$ (the green circle).
@@
  • In [herty2024multiscale], it ensures the validity of an averaging principle which reduces a singularly perturbed (or, multi–scaled) dynamics of the form (1.4), into an averaged one. This can be seen as a model reduction technique for solving multi–level optimization problems using multiple populations of interacting particles.
  • In [huang2024self], it allows to prove existence and uniqueness of an invariant probability measure $\rho_{*}^{\alpha}$, and to establish the long–time behavior of its law. Moreover, its approximation with empirical measures has been proven with an explicit rate of convergence, leading to the so-called Self–interacting CBO.
  • In [huang2024uniform], it helped to guarantee a uniform–in–time propagation of chaos for the population of particles evolving along the dynamics (1.5). This property is crucial for ensuring a stable and reliable long–term convergence, which secure the practical effectiveness of CBO methods.
The present manuscript is in the continuation of the former work [huang2024self, huang2024uniform], where the proof of global convergence is absent. Our main contributions are discussed next.
@@ 1.3. Highlights
  • Asymptotic behavior when $t\to\infty$ and $\alpha\to\infty$: We obtain the first true convergence result in the literature in the sense that both limits, $t\to\infty$ and $\alpha\to\infty$, can be taken. More specifically, we establish the asymptotic behavior of the unique invariant ($t\to\infty$) probability measure $\rho_{*}^{\alpha}$ of (1.4) as $\alpha\to\infty$. More precisely, when the function $f(\cdot)$ to be minimized has a unique global minimizer $x_{*}$, and for a given fixed parameter $\kappa$, we prove that
    $\lim_{\alpha\to\infty}\lim_{t\to\infty}\;{\mathbb{E}}\left[\,{\overline{X}}_{t}\,\right]=\lim_{\alpha\to\infty}\int_{\mathbb{R}^{d}}x\,\rho_{*}^{\alpha}(\mbox{\rm d}x)=\kappa\,\lim_{\alpha\to\infty}\mathfrak{m}_{\alpha}(\rho_{*}^{\alpha})=\kappa\,x_{*}\,,$
@@
  • Unconditional convergence: We are able to bypass classical assumptions in the literature.
@@
    • Our results hold without any prior assumptions on the initial distribution $\rho_{0}$ of (1.4) or (1.5), provided it has sufficiently many finite moments. In particular, we do not assume that a global minimizer of $f(\cdot)$ is contained within the support of $\rho_{0}$, a crucial assumption in both [carrillo2018analytical] and [fornasier2024consensus]. Very recently, [fornasier2025regularity] demonstrated that this assumption can actually be removed by instead requiring regular initial data $\rho_{0}\in H^{m}(\mathbb{R}^{d})$. In contrast, our approach does not impose any regularity conditions on $\rho_{0}$, thereby allowing cases where $\rho_{0}$ is a Dirac measure, as has been considered in the self–interacting CBO model [huang2024self].
@@
    • No regularity assumptions on the function $f(\cdot)$ are needed, besides being positive, locally Lipschitz continuous, and with a quadratic growth (see Assumption 2.1). These assumptions were used in [carrillo2018analytical] to guarantee the well-posedness of the CBO dynamics (1.3). To establish global convergence, [carrillo2018analytical, Assumption 4.1] imposes additional conditions, such as $f\in\mathcal{C}^{2}(\mathbb{R}^{d})$ and the boundedness of $\|\nabla^{2}f\|_{\infty}$. In the framework of [fornasier2024consensus, Definition 8], some additional local growth conditions on $f$ around the global minimizer are required. For our rescaled CBO we do not require these additional assumptions for the convergence proof.
@@
  • Multiple–minimizer case: Although our rescaled CBO is not specifically designed to address the multiple minimizer problem, we can still establish certain asymptotic properties in cases where the objective function $f(\cdot)$ has multiple minimizers, whether they are discrete points or compact subsets (see Theorem 3.9). Let us mention here that specially designed CBO models aimed at addressing multiple-minimizer problems have been introduced in [fornasier2025pde, bungert2022polarized].
Our results rely on two main ingredients: uniform estimates on the moments of the measure, and on a Harnack inequality on the invariant elliptic CBO equation. Indeed we need Harnack’s inequality in order to ensure that the measure has a non–null mass around minimizers, which will then allow us to prove Laplace’s principle. A summary of our findings is in section 4 at the end of the manuscript.

### v2

@@ Abstract.
We analyze the Consensus-Based Optimization (CBO) algorithm with a consensus point rescaled by a fixed parameter $\kappa\in(0,1)$. Under minimal assumptions on the objective function and the initial data, we establish its unconditional convergence to the global minimizer. Our results hold in the asymptotic regime where both the time–horizon $t\to\infty$ and the inverse–temperature $\alpha\to\infty$, providing a rigorous theoretical foundation for the algorithm’s global convergence. Furthermore, our findings extend to the case of multiple and non–discrete set of minimizers.
@@ 1. Introduction
Consensus–Based Optimization (CBO) is a derivative–free algorithm designed to tackle large–scale, highly non–convex optimization problems. This algorithm and its analytical foundations were established in [43, 13]. It has since then been extensively studied, both theoretically and practically, demonstrating its effectiveness in solving complex optimization challenges that arise in real–world applications.
Inspired by the collective behavior observed in social and biological systems, CBO exploits the interactions of multiple agents, or particles, to explore the search space in a coordinated and efficient manner. The method uses a stochastic consensus mechanism, where particles are dynamically guided toward regions of lower objective values, ultimately converging to an optimal solution. This decentralized and adaptive search strategy makes CBO particularly well-suited for problems where gradient information is unavailable or unreliable. Motivated by its broad applicability, researchers have extended and refined the original CBO framework to accommodate various settings and demonstrated strong empirical success in solving wide range of global optimization problems [23, 30, 7, 31, 14, 15, 24, 22, 26, 12, 17, 45, 6, 4, 20, 37, 19]. These advancements have further enhanced the algorithm’s robustness and efficiency, making it a versatile tool across a wide range of scientific and engineering disciplines. Typical convergence result of CBO establishes that under suitable assumptions on the objective function (e.g., continuity, isolated global minimizers, and a “localization” condition such as the so-called “well-behaved” landscape around the global minimum), the mean-field limit of CBO approximates a Dirac mass concentrated at the global minimizer when $t$ and $\alpha$ are sufficiently large. This is a probabilistic and asymptotic guarantee in the mean-field regime, not a worst-case computational complexity statement.
@@
$\text{Find }\;x_{*}\in\operatorname*{argmin}_{x\in{\mathbb{R}}^{d}}f(x),\,$
@@
where $\sigma>0$ is a real constant, $\rho_{t}^{N}:=\frac{1}{N}\sum_{i=1}^{N}\delta_{X_{t}^{i}}$ is the empirical measure associated to the particle system, and $\{B_{.}^{i}\}_{i=1}^{N}$ are $N$ independent $d$-dimensional Brownian motions. Moreover, the particle system is initialized with independent and identically distributed (i.i.d.) data $\{X_{0}^{i}\}_{i=1}^{N}$, where each $X_{0}^{i}$ is distributed according to a given measure $\rho_{0}\in\mathscr{P}({\mathbb{R}}^{d})$ for all $i\in[N]$.
Here, we employ the anisotropic diffusion in the sense that $D(X):=\mbox{diag}(|X_{1}|,\dots,|X_{d}|)$ for any $X\in{\mathbb{R}}^{d}$, which has been proven to handle high-dimensional problems more effectively [15, 24]. The current global consensus point $\mathfrak{m}_{\alpha}(\rho_{t}^{N})$ is meant to approximate the global minimum, and is defined with a weighted average of the particles as
@@
This choice of weight function $\omega_{\alpha}^{f}(\cdot)$ is motivated by the well-known Laplace’s principle [42, 21], see also [35, Appendix A.2]. Formally, this means that
(1.3) $\mathfrak{m}_{\alpha}(\rho_{t}^{N})\;\approx\;\operatorname*{argmin}_{1\leq i\leq N}f(X_{t}^{i})\quad\text{as }\alpha\to\infty.$
In other words, as $\alpha\to\infty$, the consensus point $\mathfrak{m}_{\alpha}(\rho_{t}^{N})$ converges to the position of the particle with the smallest objective value at time $t$, i.e., the best location found by the ensemble $\{X_{t}^{i}\}_{i=1}^{N}$ at time $t$. Once consensus is reached, all particles concentrate around the consensus point $\mathfrak{m}_{\alpha}(\rho_{t}^{N})$, in the sense that
(1.4) $X_{t=\infty}^{i}\;\approx\;\mathfrak{m}_{\alpha}(\rho_{t=\infty}^{N})\quad\text{for all }i\in\{1,\dots,N\}.$
Consequently, provided the number of particles $N$ is sufficiently large and $\alpha\to\infty$, the ensemble concentrates near the global minimizer $x_{*}$. We refer to [33, Section 1.2] for more details.
The convergence proof of the CBO method is typically conducted in the context of large-particle limit [29, 36]. Specifically, rather than analyzing the NN–particle system (1.1) directly, one considers the limit as $N$ approaches infinity and examines the corresponding McKean–Vlasov process $\overline{X}$$_{\cdot}$, which is governed by the following equation:
@@
$\text{with }\;\rho_{t}^{\alpha}\,=\,\text{Law}({\overline{X}}_{t}).$
@@ 1.1. The convergence problem in the literature
There are numerous results in the literature on the convergence analysis of CBO methods, which can be broadly categorized into two different approaches. For instance, in [25], the authors employ $\mathbb{E}[|\overline{X}_{t}-x_{*}|^{2}]$ as a Lyapunov functional for (1.5). Their convergence result states that for any given accuracy $\varepsilon>0$, there exists a time $T_{*}>0$ depending on $\varepsilon$, such that for $\alpha>\alpha_{0}$ (also depending on $\varepsilon$), the following holds:
@@
This approach has been further extended to establish the convergence of various CBO variants, including those incorporating general constraints [8, 3], multi-player games [18], min-max problems [9], multiple-minimizer problems [27], clustered federated learning [16], and bi-level optimization [44].
However, a limitation of this method is that it does not allow for taking the limits $t\to\infty$ or $\alpha\to\infty$, meaning that convergence is guaranteed only within a neighborhood of $x_{*}$ with accuracy $\varepsilon$. More precisely, the proof of this method relies on a quantitative version of Laplace’s principle, which requires that $\rho_{0}(B_{r_{\varepsilon}}(x^{*}))>0$. However, as $\varepsilon\to 0$, the radius $r_{\varepsilon}\to 0$, and consequently $\rho_{0}(B_{r_{\varepsilon}}(x^{*}))\to 0$, leading to a degeneracy that violates this condition.
One possible approach to achieving convergence as $t\to\infty$ is through a variance-based method described in [13]. This approach involves defining the variance function
@@
The proof is then divided into two steps. First, it is shown that under well-prepared initial data and appropriately chosen parameters, for any fixed $\alpha$, the function $V(t)$ converges to zero exponentially fast as $t\to\infty$. This implies that $\overline{X}_{t}$ converges to some consensus point $x^{\alpha}$ as $t\to\infty$. Second, it is verified that the consensus point $x^{\alpha}$ can be made arbitrarily close (up to any given accuracy $\varepsilon$) to the global minimizer $x_{*}$ by choosing some fixed $\alpha$ that is sufficiently large. This argument has been further extended to establish the convergence of such as particle swarm optimization [34], CBO constrained on the sphere [23], CBO for saddle point problems [35], and CBO with mirror descent [10]. A similar idea is also employed in the convergence proof at the particle level [40]. However, similar to the first method, this approach does not allow taking the limit $\alpha\to\infty$ and it only guarantees convergence within an $\varepsilon$–neighborhood of the global minimizer. More precisely, this method requires that the initial variance satisfies $V(0)\leq\frac{C}{\alpha}$ and that the initial distribution contains the global minimizer $x_{*}$. Consequently, as $\alpha\to\infty$, the condition $V(0)\leq\frac{C}{\alpha}$ forces the initial distribution to concentrate at $x_{*}$ already.
@@
1.2. Motivation and the rescaled CBO
As it has been observed in [32], see also [33, Remark 1.1], any Dirac measure is invariant for the dynamics (1.5) which makes it difficult to study its asymptotic properties. This non–uniqueness of an invariant probability measure is mainly due to two reasons: in the drift, the term X-X is not strong enough with respect to the mean–field term $\mathfrak{m}_{\alpha}(\cdot)$, and the diffusion degenerates. In practice, in such situations, the particle population are often observed to collapse prematurely to a Dirac delta distribution before the global minimizer has been reached (see Figure 1) due to the global minimizer not being contained in the initial distribution. Moreover, on the theoretical side, while convergence in mean-field law over finite time $t<T$ and finite $\alpha$ has been achieved in the literature (as discussed in Section 1.1), proving global convergence over an infinite time horizon has remained elusive.
Figure 1. We apply the standard CBO ($\kappa=1$ and $\delta=0$) and the rescaled CBO ($\kappa=0.01$ and $\delta=5$) particle system (1.7) to the Rastrigin function $R(x):=10+(x-1)^{2}-10\cos(2\pi(x-1))$, which has a unique global minimizer $x_{*}=1$ (the red star). The initial particles (the blue dots) are sampled uniformly in $[2,5]$ (it does not contain $x_{*}$). The simulation parameters are $N=100,\lambda=1,\sigma=0.5,\alpha=10^{15},{\rm d}t=0.01$ and $T=100$. The final output is $\mathfrak{m}_{\alpha}(\rho_{T}^{N})$ (the green circle).
Left: The standard CBO collapses prematurely to a local minimizer.
Right: The rescaled CBO finds the global minimizer.
This is the main reason behind our modification of (1.5) whose aim is to address exactly these two mentioned issues, and which has already been proposed in [33, 32, 31]. Hence, we consider a rescaled CBO given by
@@ 1.2. Motivation and the rescaled CBO
$\text{with }\;\rho^{\alpha}_{t}\,=\,\text{Law}({\overline{X}}_{t}).$
complemented with an initial condition ${\overline{X}}_{0}\sim\rho_{0}\in\mathscr{P}_{4}({\mathbb{R}}^{d})$, where $\mathds{I}_{d}$ is the $d$-dimensional identity matrix, $0<\kappa<1$ is a small positive constant, and $\delta>0$ is any positive constant. The name is justified by the fact that the contribution of the mean–field term $\mathfrak{m}_{\alpha}(\cdot)$ is re-scaled with a parameter $\kappa$ when compared to the contribution of $\overline{X}$}, thus promoting dissipation ($\overline{X}$}) over interaction ($\mathfrak{m}_{\alpha}(\cdot)$). To implement the model numerically, one considers the underlying system of $N$ interacting particles which is of the form
@@
and $\rho_{t}^{N}:=\frac{1}{N}\sum_{i=1}^{N}\delta_{X_{t}^{i}}$ is the empirical measure associated to the particle system.
Introducing the parameters $\kappa$ and $\delta$ makes all Dirac measures not invariant any more, and intuitively, creates a disconnection between the limits in the time variable $t\to\infty$ and the inverse–temperature $\alpha\to\infty$, hence allowing to easily handle the convergence of the dynamics towards the (to–be) global minimum. We refer to [33, Section 1.2] for more details.
Moreover, the positive parameter $\delta\,$ guarantees ellipticity (non–degeneracy) of the PDE satisfied by the invariant probability measure of the dynamics, and grants it nice properties. More precisely, we need $\delta>0$ in order to guarantee existence of an invariant probability measure (Theorem 3.1 below), and to ensure suitable regularity (Theorem 3.8 below). Uniqueness then follows using the $\mathbb{W}_{2}$–contraction (Theorem 3.3 below). Its value however does not influence the asymptotic analysis that we shall perform in the sequel, or the results that we aim for. Indeed, as we will later see in (4.4), the constant $\delta$ appears in a martingale term which will vanish after taking the expectation, hence its value will not influence our convergence result as long as it is positive.
To further motivate the introduction of the scaling parameter $\kappa\in(0,1)$ and the non-degenerate diffusion $\delta>0$, we present some numerical experiments in Table 1.
Choices of $\kappa$ and $\delta$ Final Output
$\kappa=1$ and $\delta=0$ The standard CBO gets trapped at a local minimizer (see Figure 1-(Left)).
$\kappa=1$ and $\delta=5$ The CBO with non-degenerate diffusion successfully locates the global minimizer.
$\kappa=0.99$ and $\delta=0$ The rescaled CBO converges to the global minimizer.
$\kappa=0.01$ and $\delta=5$ The rescaled CBO converges to the global minimizer (see Figure 1-(Right)).
Table 1. We apply the CBO dynamics (1.7) to the Rastrigin function $R(x)$ with different choices of $\kappa$ and $\delta$.
@@
  • In [31], it ensures the validity of an averaging principle which reduces a singularly perturbed (or, multi–scaled) dynamics of the form (1.6), into an averaged one. This can be seen as a model reduction technique for solving multi–level optimization problems using multiple populations of interacting particles.
  • In [32], it allows to prove existence and uniqueness of an invariant probability measure $\rho_{*}^{\alpha}$, and to establish the long–time behavior of its law. Moreover, its approximation with empirical measures has been proven with an explicit rate of convergence, leading to the so-called Self–interacting CBO.
  • In [33], it helped to guarantee a uniform–in–time propagation of chaos for the population of particles evolving along the dynamics (1.7). This property is crucial for ensuring a stable and reliable long–term convergence, which secure the practical effectiveness of CBO methods.
The present manuscript is in the continuation of the former work [32, 33], where the proof of global convergence is absent. Our main contributions are discussed next.
@@ 1.3. Highlights
  • Asymptotic behavior when $t\to\infty$ and $\alpha\to\infty$: We obtain the first true convergence result in the literature in the sense that both limits, $t\to\infty$ and $\alpha\to\infty$, can be taken. More specifically, we establish the asymptotic behavior of the unique invariant ($t\to\infty$) probability measure $\rho_{*}^{\alpha}$ of (1.6) as $\alpha\to\infty$. More precisely, when the function $f(\cdot)$ to be minimized has a unique global minimizer $x_{*}$, we prove that $\mathbb{W}_{2}(\rho^{\alpha}_{t},\rho^{\alpha}_{*})^{2}\longrightarrow 0$ when $t\to\infty$, and
    $\lim_{\alpha\to\infty}\lim_{t\to\infty}\;\mathfrak{m}_{\alpha}(\rho^{\alpha}_{t})=\lim_{\alpha\to\infty}\lim_{t\to\infty}\;\kappa^{-1}\,{\mathbb{E}}\left[\,{\overline{X}}_{t}\,\right]=x_{*}\,$,
@@
  • Unconditional global convergence: We are able to bypass classical assumptions in the literature.
@@
    • Our results hold without any prior assumptions on the initial distribution $\rho_{0}$ of (1.6) or its corresponding interacting particles system (1.7), provided it has sufficiently many finite moments. In particular, we do not assume that a global minimizer of $f(\cdot)$ is contained within the support of $\rho_{0}$, a crucial assumption in both [13] and [25]. Very recently, [28] demonstrated that this assumption can actually be removed by instead requiring regular initial data $\rho_{0}\in H^{m}(\mathbb{R}^{d})$. Or in [37], a feedback control is used to drive the particles nearby the minimizer. In contrast, our approach does not use control and does not impose any regularity conditions on $\rho_{0}$, thereby allowing cases where $\rho_{0}$ is a Dirac measure, as has been considered in the self–interacting CBO model [32].
@@
    • No regularity assumptions on the function $f(\cdot)$ are needed, besides being positive, locally Lipschitz continuous, and with a quadratic growth (see Assumption 2.1). These assumptions were used in [13] to guarantee the well-posedness of the CBO dynamics (1.5). To establish global convergence, [13, Assumption 4.1] imposes additional conditions, such as $f\in\mathcal{C}^{2}(\mathbb{R}^{d})$ and the boundedness of $\|\nabla^{2}f\|_{\infty}$. In the framework of [25, Definition 8], some additional local growth conditions on $f$ around the global minimizer are required. For our rescaled CBO we do not require these additional assumptions for the convergence proof.
@@
  • Multiple–minimizer case: Although our rescaled CBO is not specifically designed to address the multiple minimizer problem, we can still establish its asymptotic properties in the case where the objective function $f(\cdot)$ has multiple non-trivial minimizers, whether they are discrete points or compact subsets (see Theorem 4.9). More precisely, denoting $\mathcal{M}$ the set of minimizers, we prove
    $\lim\limits_{\alpha\to\infty}\,\lim\limits_{t\to\infty}\operatorname{dist}(\mathfrak{m}_{\alpha}(\rho^{\alpha}_{t}),\,\mathcal{M})=\lim\limits_{\alpha\to\infty}\,\lim\limits_{t\to\infty}\operatorname{dist}\!\left(\kappa^{-1}\,{\mathbb{E}}\left[\,{\overline{X}}_{t}\,\right],\,\mathcal{M}\right)=0.$
    Let us mention here that specially designed CBO models aimed at addressing multiple-minimizer problems have been introduced in [27, 11].
1.4. Strategy and organization
The proof of our main global convergence result can be divided into three steps as described in the following.
Step 1.
[…]
Step 2.
[…]
Step 3.
[…]
Our paper is organized as follows. We start in Section 2 by setting our main assumptions and providing some estimates, in particular on the moment bound. Then Section 3 shall be devoted to the invariant measure. Therein, we shall prove its existence, the long-time limit ($t\to\infty$), and other of its important properties: uniform-in-$\alpha$ second moment and Harnack’s inequality. Subsequently, the limit $\alpha\to\infty$ will be tackled in Section 4. In particular, we shall prove a Laplace’s principle, then the global convergence towards the minimizer. An analogous convergence result will also be provided in the case of multiple non-trivial minimizers. Section 5 contains some numerical experiments and discussion regarding our model and results. Finally, a summary of our findings is in Section 6 at the end of the manuscript, together with future research directions. Our paper concludes with Appendix A which contains detailed computations for verifying Theorem 3.1.